OpenStudy (anonymous):

Solve. x+4y+z=-10 3x-3y+6z=-21 x+2y+2z=-10

OpenStudy (anonymous):

do you know how to convert this to matrix and solve with matrix algebra?

OpenStudy (anonymous):

I ave to use substitution and elimination .

OpenStudy (anonymous):

have*

OpenStudy (anonymous):

google those then :) i was thinking substitution but that will take a while.

OpenStudy (anonymous):

how do i do it ?

OpenStudy (anonymous):

I'll give you definitions. Subs method is solving for one variable in one equation and plugging into the other equations.

OpenStudy (anonymous):

http://www.purplemath.com/modules/systlin5.htm Here is eliminations I guess they multiply to eliminate constants and variable.s

OpenStudy (inkyvoyd):

HINT: first and third equations give you opportunity to easily eliminate x

OpenStudy (inkyvoyd):

HINT: then use the first and second equations to eliminate x again. The resulting two equations with x eliminated can both be solved for y and z. Then, easy as usual, you can just plug y and z back in to get x.

OpenStudy (inkyvoyd):

Try doing it a bit yourself, and I will help you if you need further help :)

OpenStudy (anonymous):

So elimination? and then which one do i solve for?

OpenStudy (inkyvoyd):

Okay. When you see this kind of system, the first goal is to simplify it. We do this by eliminating variables. That makes sense right?

OpenStudy (inkyvoyd):

x+4y+z=-10 call this equation 1 (I will use e1 for short) 3x-3y+6z=-21 call this equation 2 (I will use e2 for short) x+2y+2z=-10 call this equation 3 (I will use e3)

OpenStudy (inkyvoyd):

by combining different equations, we can get simplified results.

OpenStudy (anonymous):

which ones should i combine first ?

OpenStudy (inkyvoyd):

Btw, combining does not mean adding. You can combine any multiple of the entire equation (remember that means negatives included)

OpenStudy (anonymous):

huuhhhhhhh ?

OpenStudy (inkyvoyd):

Okay. I'll worth through a little bit of this problem, but you might have to give me 5 mins, cause I have to do some stuff and I'm also helping someone else with long division.

OpenStudy (inkyvoyd):

Subtract equation 3 from equation 1. In other words, x+4y+z-(x+2y+2z)=-10-(-10) Does this make sense?

OpenStudy (inkyvoyd):

Okay. What didn't make sense?

OpenStudy (anonymous):

subtracting it ?

OpenStudy (inkyvoyd):

Okay. I'll do it another way. x+4y+z=-10 is e1 (equation one) right?

OpenStudy (inkyvoyd):

x+2y+2z=-10 is equation 3 right?

OpenStudy (anonymous):

is the answer .. (1,-2,-3) ?

OpenStudy (inkyvoyd):

How'd you get that :)

OpenStudy (anonymous):

I took the answers and put them in until i got -10 = -10 hahaha

OpenStudy (anonymous):

x = -4, y = -1, z = -2

OpenStudy (anonymous):

is that correct??

OpenStudy (anonymous):

mines is :P

OpenStudy (anonymous):

plug them in and see if they work.